OpenStudy (maheshmeghwal9):

Solve:-- (1/x)+(1/50-x)=1/12

OpenStudy (anonymous):

this a long way from proving a function is continuous at a point,isn't it?

OpenStudy (maheshmeghwal9):

ya ,

OpenStudy (anonymous):

add on the left and then cross multiply

OpenStudy (maheshmeghwal9):

k!

OpenStudy (anonymous):

dont forget about the lcd

OpenStudy (anonymous):

\[\frac{1}{x}+\frac{1}{50-x}\] \[\frac{50-x+x}{x(50-x)}\] \[\frac{50}{x(50-x)}=\frac{1}{2}\] \[100=x(50-x)\] etc

OpenStudy (maheshmeghwal9):

Thanx!

Parth (parthkohli):

Remember, \(\Large \color{purple}{\rightarrow {a \over b} + {c \over d} = {ad + cb \over bd } }\)

OpenStudy (maheshmeghwal9):

K! dude :)

OpenStudy (maheshmeghwal9):

thanx!